1.1

> f1 := plot3d(x^2 + y^2 - 5, x=-3..5, y= -Pi..Pi, view = -5..3,
color=blue):
> with (plots):
> f2 := plot3d(sin(x)+sin(y), x=-3..5, y= -Pi..Pi, view = -3..3,
color=green):
> display(f1, f2);

1.1.1 Exercises:

11) Plot the functionsz = x^{2} + 2 y^{2} andz = 2 x − 3 + 4 y on the same axes. Plot over the rectangle with x and y between -1 and 3.

12) Regraph the same pair of functions, this time constraining the z-range between -1 and 7.

These 2 surfaces have a simple relationship. Describe the relationship and replot the graph viewed at an angle to make the relationship clear visually.

The relationship is....... The viewing angle that makes this obvious is ......